OpenStudy (anonymous):

When differentiating implicitly, is d/dx [2x] = 0 or 2x*d/dx? Here's the original equation: 2x - 3xy^2 = 10

OpenStudy (luigi0210):

Are we differentiating twice?

OpenStudy (anonymous):

Uhh I don't think so. I'm trying to find the slope of this curve when y=-2. So I first set this equation as d/dx [2x - 3xy^2] = d/dx [10] and made that simpler by d/dx [2x] + d/dx [-3xy^2] = 0. But I'm not sure what d/dx [2x] would result in.... please help!!

OpenStudy (luigi0210):

Oh, since it's in respect to x, it would just be 2

OpenStudy (anonymous):

Okay, thanks very much! :)

OpenStudy (luigi0210):

You're welcome :)